Grade Select GradeTrigonometry

y = 2sin¯13x

Hassi Mohammed
11 Years agoGrade Select Grade

1 Answer

Latika Leekha
11 Years ago
Hello student,
As u haven’t mentioned what exactly is the question, I am assuming that you need the value of x.
We have y = 2sin-1 3x
So, y/2 = sin -1 3x
Hence, x = (1/3) sin (y/2).